OpenStudy (anonymous):

1-cos^2(theta)/1+cos^2(theta)+2cos^2(theta) simplify

OpenStudy (anonymous):

\[\left(\frac{1-\cot^2\left(\theta\right)}{1+\cot^2\left(\theta\right)}\right)+2\cos^2\left(\theta\right)\]

OpenStudy (anonymous):

\[1 + \cot^2(x) = cosec^2(x)\]

OpenStudy (anonymous):

\[\frac{1- \cot^2(x)}{cosec^2(x)} + \cos^2(x) \implies \frac{1}{cosec^2(x)} - \frac{\cot^2(x)}{cosec^2(x)} + \cos^2(x)\]

OpenStudy (anonymous):

\(\large \frac{1}{cosec^2(x)}\) = ??

OpenStudy (anonymous):

so simplified it just equals 1/ cosec^2 (x)

OpenStudy (anonymous):

??

OpenStudy (anonymous):

dont have to solve jsut simplify

OpenStudy (anonymous):

I ask you something.. Answer that..

OpenStudy (anonymous):

Simplifying means we need to solve it and simplify more, in its simplest form we have to represent it..

OpenStudy (anonymous):

1/cosec = sin

OpenStudy (anonymous):

so 1 / csc^2 = sin^2?

OpenStudy (anonymous):

Yes, so: \[\sin^2(x) - \cot^2(x) \cdot \sin^2(x) + \cos^2(x)\]

OpenStudy (anonymous):

Now, \(sin^2(x) + cos^2(x)\) = ??

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

Yep: it becomes :\[\implies 1 - \cot^2(x) \cdot \sin^2(x)\]

OpenStudy (anonymous):

Now, do you know any formula for \(cot(x)\) in terms of \(sin(x)\) and \(cos(x)\) ??

OpenStudy (anonymous):

1/tan = cot

OpenStudy (anonymous):

so 1/sin/cos

OpenStudy (anonymous):

or simply cos/sin Okay??

OpenStudy (anonymous):

ah yea simplfied you right

OpenStudy (anonymous):

\[\cot(x) = \frac{\cos(x)}{\sin(x)}\]

OpenStudy (anonymous):

Put it there..

OpenStudy (anonymous):

\[\implies 1 - \frac{\cos^2(x)}{\sin^2(x)} \cdot \sin^2(x)\]

OpenStudy (anonymous):

\[\implies 1 - \frac{\cos^2(x)}{\cancel{\sin^2(x)}} \cdot \cancel{\sin^2(x)} \implies 1 - \cos^2(x) = ??\]

OpenStudy (anonymous):

sin^2(x)

OpenStudy (anonymous):

yep.. Good..!!! :)

OpenStudy (anonymous):

you're a lifesaver!

OpenStudy (anonymous):

Really?? But OS is giving title of "Champion" to me... :P Just kidding..!! Actually, I like Trigonometry, it is my favorite topic from MATHS.. :)

OpenStudy (anonymous):

Haha! Ya math is kinda fuzzy for me taht's my im chosing to major in chemistry! lol

OpenStudy (anonymous):

oooo

OpenStudy (anonymous):

what about that 2cos^2(x) at end